# and the Quantum Hole Argument

### v1

@@ Abstract
The study of quantum reference frames (QRFs) is motivated by the idea of taking into account the quantum properties of the reference frames that we use, explicitly or implicitly, in our description of physical systems. Like a classical reference frame, a QRF can be used to define physical quantities such as time, position, momentum, and spin relationally. Unlike its classical analogue, it relativises the notions of superposition and entanglement of quantum systems. Here, we provide a novel explanation for the frame-dependence of superposition and entanglement by tracing it back to the question of how configurations or locations are identified across different branches in superposition. We show that, in the presence of symmetries, whether a system is in “the same” or “different” configurations across the branches depends on the choice of QRF. Thus, sameness and difference — and, as a result, superposition and entanglement — lose their absolute meaning. We apply these ideas to the context of semi-classical spacetimes in superposition and use coincidences of four scalar fields to construct a comparison map between the spacetime points in the different branches. This allows us to determine whether a given event is located at “the same” or “different” points in the superposed spacetimes. Because this feature depends on the choice of QRF, we argue that the localisation of an event should not be seen as an inherent property of the event. This alleviates previously voiced concerns that QRF changes could have empirical consequences for interference experiments, such as the Bose et al.-Marletto-Vedral proposal. Moreover, it implies that the number of events is equal in both the flat and the curved spacetime implementations of quantum-controlled causal order. We conclude with the “quantum hole argument” as a generalisation for the quantum context of Einstein’s famous hole argument, arguing that not just spacetime points but also their identification and the localisation of events across manifolds in superposition lose their absolute physical meaning in the presence of quantum symmetries.
@@ I Introduction
When describing a physical system, it is very common to do so with respect to a reference frame — a ruler used to determine the position of a particle, for example, or a clock, which tracks the time that elapses while it is moving. Furthermore, one can take the view that “the particle’s position” or “the tracking of time” are meaningful only in relation to the degrees of freedom of the system that serves as the reference frame or clock. Such a relational account of physical systems and dynamics seems to be a fundamental aspect of how we describe the world. Usually, the reference systems that we employ, either explicitly or implicitly, are treated as purely classical objects with well-defined properties. But what happens if we take into account the possibility that the reference system itself exhibits quantum properties? This question has motivated a recent wave of research on quantum reference frames (QRFs) [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23], investigating what the world looks like with respect to a quantum system. In particular, it has been found that the quantum properties of the frame, such as superposition or entanglement, get transferred onto the relational description of the surrounding systems. As a result, a QRF transformation can change whether a system appears to be in superposition of different configurations, and whether it is entangled with another system, thus rendering these features frame-dependent [2, 8]. If we assume that the covariance of physical laws under classical changes of reference frame extends linearly to encompass changes of QRF [2], then the latter should not affect the physical situation. Thus, whether a system is in superposition or is entangled, and how these properties change dynamically, becomes a mere matter of perspective.
@@
The frame-dependence of quantum properties is already well understood at the formal level, within several different frameworks for QRFs [24, 1, 2, 8, 3, 7, 14, 13, 22, 23]. However, the meaning and usage of QRFs is still not fully clarified on a conceptual level. The purpose of the present work is to abstract from the mathematical details of the various formalisms and applications and to crystallise their common core by embedding several recent works on QRFs into a single unified picture. It provides a visualisation of different choices of QRFs and how they are connected. This, we believe, can greatly enhance the intuitive understanding of QRFs and allow for a more pedagogical introduction into the subject. With the help of concrete toy examples, we offer a new understanding of the frame-dependence of superposition and entanglement under QRF changes: Ultimately, whether a system is in “the same” or “different” configurations across the branches — and thereby, whether it is in a superposition or not — depends on which QRF we use to identify and compare these configurations. These notions of identification and comparison change under a QRF transformation. As a result, sameness and difference lose their absoluteness.
@@
In addition to providing a more intuitive access to QRFs from a purely quantum mechanical perspective, the conceptual nature of the framework also allows for its application to general relativistic scenarios. This enables us to explore the conceptual implications of an invariance under QRF transformations for the diffeomorphism group in the context of semi-classical spacetimes in superposition (as considered, e.g., in [33, 34, 17, 16, 35, 36, 18, 37, 38, 19]). This regime is of particular interest as it might allow us to better understand spacetime at the quantum level, similar to how the classical diffeomorphism invariance of general relativity has significantly illuminated our classical understanding of spacetime. We find that, in this context, a choice of QRF comes with a notion of how to identify spacetime points across the superposition. Note that a priori there does not exist any preferred way of comparing or identifying the locations of objects across different spacetimes in superposition. There have been some attempts in the literature at constructing such a “threading” of points between the different manifolds, on the classical [39] and the quantum level [40, 41]. In [33, 34], the authors take a first step towards implementing such a threading in the context of QRFs by using the location of a probe particle to identify some points of the different manifolds in superposition. Nevertheless, in a diffeomorphism-invariant theory, the position of a single particle is not sufficient to specify how to identify points across entire spacetimes in superposition. Here, we extend this idea by replacing the probe particle by four scalar fields and using their coincidences to construct a comparison map between all spacetime points in the superposition. This provides us with the necessary structure to define whether an event is located at “the same” or “different” points across the spacetimes in superposition — whether it is localised or delocalised. While we show that the localisation of events depends on the choice of QRF, we argue that, nonetheless, QRF transformations do not have empirical consequences for phenomena such as interference and recombination. Moreover, the insight that localisation is frame-dependent helps to shed light on the notion of a “quantum event” and in particular the question of how to count events in the context of superpositions of spacetimes, which also plays an important role in the study of indefinite causal order [42, 43, 44, 45, 46, 47, 19]. Furthermore, the understanding of QRFs as providing a preferred identification of points across spacetimes in superposition has interesting implications for a possible quantum generalisation of the hole argument, which we discuss at the end of this article.
@@ II The Space of Models and Representational Conventions
In the case of general relativity, one orbit would contain all diffeomorphically equivalent spacetimes; that is, including the metric and matter fields. Because $G$ is a symmetry group of the theory, we say that models in the same orbit are physically equivalent or indistinguishable. Physically distinct models, such as $\varphi$ and $\varphi^{\prime}$, are in different orbits. As a consequence of the symmetry of our theory, the space of models contains a lot of redundancy — to obtain a complete description of all physically distinct states of the system our theory is describing, it would suffice to know one representative of each orbit. In gauge theories, one usually deals with this redundancy by fixing a gauge. In generally covariant theories, one fixes a coordinate system. What is often implicitly assumed in these procedures is that there is a unique prescription, such as the Coulomb gauge condition $\vec{\nabla}\cdot\vec{A}=0$ in electrodynamics, which applies to all the models equally. [^2] In the following, we are mainly going to be referring to gauge symmetries and gauge fixings rather than general covariance and coordinate choices. However, our framework applies to generally covariant theories as well, as we will see in Sec. V. That is, the same gauge condition is imposed on all the different orbits $\mathcal{O}_{\varphi}$ to pick out a preferred representative $\varphi^{\ast}\in\mathcal{O}_{\varphi}$. This, however, is not necessary; indeed, in theories more complicated than electrodynamics, it is not even possible. [^3] The Gribov obstruction [48] shows that no single such prescription can cover all of the orbits. The symmetry of the theory grants a much larger freedom in the choice of representatives. In principle, we can choose arbitrary prescriptions, that may pick out different representatives in the different orbits. In other words, we can use different gauge fixing prescriptions for different possible states of the world. This crucial insight also lies behind the QRF formalism, but we will discuss this in more detail below. The most general way of removing the redundancy in our description is to simply pick one unique representative for each orbit. This can be achieved via an injective map
@@
where $\left[\Phi\right]$ denotes the space of equivalence classes $\left[\varphi\right]$ with respect to the equivalence relation $\sim$, which is defined by $\varphi\sim\varphi^{\prime}$ if and only if there exists a $g$ such that $\varphi^{\prime}=\varphi^{g}$. Depending on the theory, one might want the map to respect certain structures, such as smoothness or differentiability on the space of models. In [32], the map $\sigma$ is called a representational convention. For gauge theories with a space of models which has a principal fibre bundle structure, it would instantiate a section or a slice (see e.g. [49, 50]) through the fibre bundle.
@@
Figure 3: Given a representational convention $\sigma$, we can define the corresponding counterpart relation $\mathrm{Counter}_{\sigma}(\varphi,\varphi^{\prime})$ that prescribes how to compare models $\varphi$ and $\varphi^{\prime}$. In particular, it aligns the two models by lowering $\varphi$ to the section $\sigma$ using $g_{\sigma}(\varphi)$ and then raising it again with $g_{\sigma}(\varphi^{\prime})^{-1}$. In other words, it puts the two models on equal footing, allowing a direct comparison between different possible configurations.
@@
In order to see this and to make the connection to QRFs alluded to above, let us now make the crucial step that takes the notions introduced so far to the quantum level: rather than considering a set of models in $\Phi$ as a collection of possible configurations (states of our system), we take these models to denote the different branches of a quantum state of the system, i.e. branches or summands of a superposition. Note that, in this way, the symmetry group $G$ induces a preferred basis for models, in terms of which superpositions of configurations are expressed. The physical requirement, so as to distinguish such a superposition from a classical mixture, is to measure a quantity whose eigenbasis does not commute with the preferred one. In the present paper, we do not explicitly model such other bases. However, in the translation group example, described in the next Section, where the preferred basis is the position basis, one could straightforwardly express the states and observables in these other bases as well. In general, it might be challenging to perform a measurement in the conjugate basis, in particular in the context of models involving spacetime.
@@
Keeping these considerations in mind, let us now take the system’s total state to be in a superposition of models $\varphi$ and $\varphi^{\prime}$ (see Fig. 1) rather than treating $\varphi$ and $\varphi^{\prime}$ as different yet unrelated possible states of our system. Importantly, we do not grant any physical meaning to superpositions of models which live on the same orbit and instead view them as equivalent to any single representative of the equivalence class. The orbit-wise freedom in the choice of representational convention then translates to a branch-wise fixing of the gauge in the superposition of configurations. [^5] Note that in the following we will assume that the gauge-fixing section is chosen to be smooth across orbits. Moreover, a change from one representational convention to another is now implemented by a branch-dependent choice of which group element to act with — just like a QRF transformation. [^6] The idea of gauge transforming “each individual component of the superposition independently” [51, p. 2] also underlies the complete quantum gauge transformations of Rovelli [51]. This strongly suggests the connection between representational conventions for configurations in superposition and QRFs: A specific choice of representational convention $\sigma$ corresponds to the choice of a QRF relative to which the state of the system is described. This is illustrated in several concrete examples further below.
@@ III The Translation Group: Massive Object in Superposition
Now consider a superposition of different models. A change from one representational convention $\sigma$ in the space of models to another $\tilde{\sigma}$ can now be seen as a change from one position reference frame (one stipulation of which point is the spatial origin) to another. A model-independent transformation, such as a standard classical reference frame transformation, would shift the origin by the same amount in each branch of the superposition. That corresponds to applying the same group element of the translation group to each model in the total configuration (see Fig. 4(a)). A model-dependent transformation, such as a quantum reference frame transformation, on the other hand, acts with a different group element, i.e. translation, in each branch (see Fig. 4(b)). This is commonly referred to as a quantum-controlled transformation.
(a) Model-independent, classical translation
(b) Model-dependent, quantum-controlled translation
Figure 4: Graphical illustration of the difference between model-independent and model-dependent translations. Classical translations are commonly described as model-independent transformations, which shift the models by a constant amount through applying the same translation $\mathbb{T}_{a}$ in each branch of the superposition. Quantum translations, on the other hand, will, in general, apply different translations $\mathbb{T}_{a}$ and $\mathbb{T}_{a^{\prime}}$ in the different branches. As before, the dotted lines depict two sections associated with different representational conventions.
@@
Figure 6: The space of models for a system consisting of a massive object $M$ (illustrated by the Earth) and a particle $P$ (illustrated by the atom). This figure depicts a superposition of two distinct physical configurations — $M$ being to the left or to the right of $P$ — and two different representational conventions $\sigma$ and $\tilde{\sigma}$. The straight line represents the section $\sigma$ in which the zero-point of the coordinate system is associated with the position of $P$, marked by the flag. In this reference frame, $M$ is at different positions across the branches. The tilted line represents the section $\tilde{\sigma}$ in which the zero-point is associated with the position of $M$, marked again by the flag, while $P$ is in a superposition of positions. The change from one representational convention to another is implemented by a quantum translation. The images outside the space of models illustrate the frame-dependence of the superposition.
@@
relative to the particle. Note that the state of the particle relative to its own reference frame factorises out. This is a general feature of states given relative to QRFs: the relative state of the frame itself factorises out [8] and is therefore sometimes omitted [2]. Now, if the two-particle system is described, for example, by Newtonian gravity, the overall configuration of the two systems is translation-invariant. For simplicity, we exclude the rotational symmetry of the theory and consider only the invariance under translations in one dimension. As a consequence, any two models that are related by a rigid translation live on the same orbit. This is the case for models with the same relative distance $d_{MP}=x_{M}-x_{P}$. Thus in this example, each orbit can be labelled by the relative distance between the two systems $P$ and $M$. As the (signed) relative distance between $M$ and $P$ differs across the branches in our setup, we are dealing with a superposition of two models that are physically distinguishable and therefore live on different orbits in the space of models (see Fig. 6). [^9] Note that we would consider superpositions of models living on the same orbit as “false” superpositions. In the present context, this would correspond to a superposition of two configurations with the same relative distance between $M$ and $P$ but distinct centre of mass position.
@@
Consider as an example a three-particle system, whose positions are denoted by $q_{1},q_{2},q_{3}$. When applying a standard, model-independent translation, the position of each system is shifted by the same fixed amount: $q_{i}\to q_{i}+a$, $i=1,2,3$. For the model-dependent translation, on the other hand, the amount by which the positions are shifted can depend on the model $\varphi=(q_{1},q_{2},q_{3})$ and thus on the configuration of the other particles. This is the case, for example, when changing from the reference frame of Particle 1 into the reference frame of Particle 2, even at the classical level. In the former, what we mean by “the position of Particle 3” is the “the position of Particle 3 relative to Particle 1”, that is $Q_{3}^{(1)}=q_{3}-q_{1}$. Upon changing into the reference frame of Particle 2, on the other hand, “the position of Particle 3” is now $Q_{3}^{(2)}=Q_{3}^{(1)}+q_{1}-q_{2}=q_{3}-q_{2}$, namely “the position of Particle 3 relative to Particle 2”. In this sense, this transformation “mixes” the degrees of freedom of the different subsystems and thereby changes the notion of subsystem. While this change of subsystem structure is a well-understood feature at the classical level, it develops new implications at the quantum level, where it is reflected in a refactorisation of the Hilbert space and thereby the frame-dependence of entanglement.

### v2

@@ Abstract
The study of quantum reference frames (QRFs) is motivated by the idea of taking into account the quantum properties of the reference frames used, explicitly or implicitly, in our description of physical systems. Like classical reference frames, QRFs can be used to define physical quantities relationally. Unlike their classical analogue, they relativise the notions of superposition and entanglement. Here, we explain this feature by examining how configurations or locations are identified across different branches in superposition. We show that, in the presence of symmetries, whether a system is in “the same” or “different” configurations across the branches depends on the choice of QRF. Hence, sameness and difference — and thus superposition and entanglement — lose their absolute meaning. We apply these ideas to the context of semi-classical spacetimes in superposition and use coincidences of four scalar fields to construct a comparison map between spacetime points in the different branches. This reveals that the localisation of an event is frame-dependent. We discuss the implications for indefinite causal order and the locality of interaction and conclude with a generalisation of Einstein’s hole argument to the quantum context.
@@ I Introduction
When describing a physical system, it is very common to do so with respect to a reference frame — a ruler used to determine the position of a particle, for example, or a clock, which tracks the time that elapses while it is moving. Furthermore, one can take the view that “the particle’s position” or “the tracking of time” are meaningful only in relation to the degrees of freedom of the system that serves as the reference frame or clock. Such a relational account of physical systems and dynamics seems to be a fundamental aspect of how we describe the world. Usually, the reference systems that we employ, either explicitly or implicitly, are treated as purely classical objects with well-defined properties. But what happens if we take into account the possibility that the reference system itself exhibits quantum properties? This question has motivated a recent wave of research on quantum reference frames (QRFs) [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23], investigating what the world looks like with respect to a quantum system. In particular, it has been found that the quantum properties of the frame, such as superposition or entanglement, get transferred onto the relational description of the surrounding systems. As a result, a QRF transformation can change whether a system appears to be in superposition of different configurations, and whether it is entangled with another system, thus rendering these features frame-dependent [2, 8]. If we assume that the covariance of physical laws under classical changes of reference frame extends linearly to encompass changes of QRF [2], then the latter should not affect the physical situation. Thus, whether a system is in a superposed or entangled state, and how these properties change dynamically, becomes a mere matter of perspective.
@@
The frame-dependence of quantum properties is already well understood at the formal level, within several different frameworks for QRFs [24, 1, 2, 8, 3, 7, 14, 13, 22, 23]. However, the meaning and usage of QRFs is still not fully clarified on a conceptual level. The purpose of the present work is to abstract from the mathematical details of a subset of the various formalisms and applications and to crystallise their common core by embedding several recent works on QRFs [2, 5, 8, 11, 17, 18, 20, 19] into a single unified picture. It provides a visualisation of different choices of QRFs and how they are connected. This, we believe, can greatly enhance the intuitive understanding of QRFs and allow for a more pedagogical introduction into the subject. With the help of concrete toy examples, we offer a new understanding of the frame-dependence of superposition and entanglement under QRF changes: Ultimately, whether a system is in “the same” or “different” configurations across the branches — and thereby, whether it is in a state of superposition or not — depends on which QRF we use to identify and compare these configurations. These notions of identification and comparison change under a QRF transformation. As a result, sameness and difference lose their absoluteness.
@@
In addition to providing a more intuitive access to QRFs from a purely quantum mechanical perspective, the conceptual nature of the framework also allows for its application to general relativistic scenarios. This enables us to explore the conceptual implications of an invariance under QRF transformations for the diffeomorphism group in the context of semi-classical spacetimes in superposition (as considered, e.g., in [33, 34, 17, 16, 35, 36, 18, 37, 38, 19]). [^1] Note that when referring to “(semi-classical) spacetimes in superposition”, we mean that the metric is in a superposition of states, each peaked around a classical configuration of the metric. More generally, in a slight abuse of language, we will often refer to a system/spacetime/manifold/etc.“in superposition” — this is a short form for the quantum state of the system/spacetime/manifold/etc. being in superposition. As was stressed by Hardy in his work on the quantum equivalence principle [39], this regime is of particular interest as it might allow us to better understand spacetime at the quantum level, similar to how the classical diffeomorphism invariance of general relativity has significantly illuminated our classical understanding of spacetime. Note that we do not want to claim that we solve the many mathematical difficulties in dealing with the diffeomorphism group and general spacetime metrics, in particular the problems of defining well-behaved measures. Rather, we focus on what we can say by making a few core assumptions while these problems remain open. We find that, in this context, a choice of QRF comes with a notion of how to identify spacetime points across the superposition. Note that a priori there does not exist any preferred way of comparing or identifying the locations of objects across different spacetimes in superposition. There have been some attempts in the literature at constructing such a “threading” of points between the different manifolds, on the classical [40] and the quantum level [39, 41]. In [33, 34], the authors take a first step towards implementing such a threading in the context of QRFs by using the location of a probe particle to identify some points of the different manifolds in superposition. Nevertheless, in a diffeomorphism-invariant theory, the position of a single particle is not sufficient to specify how to identify points across entire spacetimes in superposition. Here, we extend this idea by replacing the probe particle by four scalar fields and using their coincidences to construct a comparison map between all spacetime points in the superposition. This provides us with the necessary structure to define whether an event is located at “the same” or “different” points across the spacetimes in superposition — whether it is localised or delocalised. While we show that the localisation of events depends on the choice of QRF, we argue that, nonetheless, QRF transformations do not have empirical consequences for phenomena such as interference and recombination. Moreover, the insight that localisation is frame-dependent helps to shed light on the notion of a “quantum event” and in particular the question of how to count events in the context of superpositions of spacetimes, which also plays an important role in the study of indefinite causal order [42, 43, 44, 45, 46, 47, 19]. Furthermore, the understanding of QRFs as providing a preferred identification of points across spacetimes in superposition has interesting implications for a possible quantum generalisation of the hole argument, which we discuss at the end of this article.
@@ II The Space of Models and Representational Conventions
In the case of general relativity, one orbit would contain all diffeomorphically equivalent spacetimes; that is, including the metric and matter fields. Because $G$ is a symmetry group of the theory, we say that models in the same orbit are physically equivalent or indistinguishable. Physically distinct models, such as $\varphi$ and $\varphi^{\prime}$, are in different orbits. As a consequence of the symmetry of our theory, the space of models contains a lot of redundancy — to obtain a complete description of all physically distinct states of the system our theory is describing, it would suffice to know one representative of each orbit. In gauge theories, one usually deals with this redundancy by fixing a gauge. In generally covariant theories, one fixes local coordinate systems. What is often implicitly assumed in these procedures is that there is a unique prescription, such as the Coulomb gauge condition $\vec{\nabla}\cdot\vec{A}=0$ in electrodynamics, which applies to all the models equally. [^3] In the following, we are mainly going to be referring to gauge symmetries and gauge fixings rather than general covariance and coordinate choices. However, our framework applies to generally covariant theories as well, as we will see in Sec. V. That is, the same gauge condition is imposed on all the different orbits $\mathcal{O}_{\varphi}$ to pick out a preferred representative $\varphi^{\ast}\in\mathcal{O}_{\varphi}$. This, however, is not necessary; indeed, in theories more complicated than electrodynamics, it is not even possible. [^4] The Gribov obstruction [48] shows that no single such prescription can cover all of the orbits. Note that it has also been explored in the context of QRFs in [4, 49, 50]. The symmetry of the theory grants a much larger freedom in the choice of representatives. In principle, we can let the way we choose a representative vary arbitrarily between the different orbits. In other words, we can use different gauge fixing prescriptions for different possible states of the world. This crucial insight also lies behind the QRF formalism, but we will discuss this in more detail below. The most general way of removing the redundancy in our description is to simply pick one unique representative for each orbit. This can be achieved via an injective map
@@
where $\left[\Phi\right]$ denotes the space of equivalence classes $\left[\varphi\right]$ with respect to the equivalence relation $\sim$, which is defined by $\varphi\sim\varphi^{\prime}$ if and only if there exists a $g$ such that $\varphi^{\prime}=\varphi^{g}$. In [32], the idea of this injective map $\sigma$ — which, in general, does not necessarily respect all structures on $\Phi$, such as being differentiable — is called a representational convention. But of course, one can require such structures to be preserved. For gauge theories, with a space of models which has a principal fibre bundle structure, $\sigma$ would be a (not in general global) section or slice (see e.g. [51, 52]) through the fibre bundle.
@@
Figure 3: The counterpart relation. Given a representational convention $\sigma$, we can define the corresponding counterpart relation $\mathrm{Counter}_{\sigma}(\varphi,\varphi^{\prime})$ that prescribes how to compare models $\varphi$ and $\varphi^{\prime}$. In particular, it aligns the two models by lowering $\varphi$ to the section $\sigma$ using $g_{\sigma}(\varphi)$ and then raising it again with $g_{\sigma}(\varphi^{\prime})^{-1}$. In other words, it puts the two models on equal footing, allowing a direct comparison between different possible configurations.
@@
In order to see this and to make the connection to QRFs alluded to above, let us now make the crucial step that takes the notions introduced so far to the quantum level: rather than considering a set of models in $\Phi$ as a collection of possible configurations (states of our system), we take these models to denote the different branches of a quantum state of the system, i.e. branches or summands of a superposition. Note that, in this way, the symmetry group $G$ induces a preferred basis for models, in terms of which superpositions of configurations are expressed. The physical requirement, so as to distinguish such a superposition from a classical mixture, is to measure a quantity whose eigenbasis does not commute with the preferred one. In the present paper, we do not explicitly model such other bases. But of course one can express states and observables in an arbitrary basis of the Hilbert space. [^6] Here, we set aside subtleties that may arise in the context of field theoretic models and in particular in the context of spacetimes in superposition that will be discussed in Sec. V. In particular, in the translation group example described in the next section, the preferred basis is the position basis but one could straightforwardly express the states and observables in other bases as well. In general, it might be challenging to perform a measurement in the conjugate basis, in particular in the context of models involving spacetime.
@@
Keeping these considerations in mind, let us now take the system’s total state to be in a superposition of models $\varphi$ and $\varphi^{\prime}$ (see Fig. 1) rather than treating $\varphi$ and $\varphi^{\prime}$ as different yet unrelated possible states of our system. Importantly, we do not grant any physical meaning to superpositions of models which live on the same orbit and instead view them as equivalent to any single representative of the equivalence class. The orbit-wise freedom in the choice of representational convention then translates to a branch-wise fixing of the gauge in the superposition of configurations. [^7] Note that in the following we will assume that the gauge-fixing section is chosen to be smooth across orbits. Moreover, a change from one representational convention to another is now implemented by a branch-dependent choice of which group element to act with — just like a QRF transformation. [^8] The idea of gauge transforming “each individual component of the superposition independently” [53, p. 2] also underlies the complete quantum gauge transformations of Rovelli [53]. This strongly suggests the connection between representational conventions for configurations in superposition and QRFs: A specific choice of representational convention $\sigma$ corresponds to the choice of a QRF relative to which the state of the system is described. This is illustrated in several concrete examples further below.
@@ III The Translation Group: Massive Object in Superposition
Now consider a superposition of different models. A change from one representational convention $\sigma$ in the space of models to another $\tilde{\sigma}$ can now be seen as a change from one position reference frame (one stipulation of which point is the spatial origin) to another. A model-independent transformation, such as a standard classical reference frame transformation, would shift the origin by the same amount in each branch of the superposition. That corresponds to applying the same group element of the translation group to each model in the total configuration (see Fig. 4). A model-dependent transformation, such as a quantum reference frame transformation, on the other hand, acts with a different group element, i.e. translation, in each branch (see Fig. 4). This is commonly referred to as a quantum-controlled transformation.
Figure 4: Graphical illustration of the difference between model-independent and model-dependent translations. (a) Classical translations are commonly described as model-independent transformations, which shift the models by a constant amount through applying the same translation $\mathbb{T}_{a}$ in each branch of the superposition. (b) Quantum translations, on the other hand, will, in general, apply different translations $\mathbb{T}_{a}$ and $\mathbb{T}_{a^{\prime}}$ in the different branches. As before, the dotted lines depict two sections associated with different representational conventions.
@@
Figure 6: The space of models for a system consisting of a massive object $M$ and a particle $P$. This figure depicts a superposition of two distinct physical configurations — $M$ being to the left or to the right of $P$ — and two different representational conventions $\sigma$ and $\tilde{\sigma}$. The straight line represents the section $\sigma$ in which the zero-point of the coordinate system is associated with the position of $P$, marked by the flag. In this reference frame, $M$ is at different positions across the branches. The tilted line represents the section $\tilde{\sigma}$ in which the zero-point is associated with the position of $M$, marked again by the flag, while $P$ is in a superposition of positions. The change from one representational convention to another is implemented by a quantum translation. The images outside the space of models illustrate the frame-dependence of the superposition.
@@
relative to the particle. Note that the state of the particle relative to its own reference frame factorises out. This is a general feature of states given relative to QRFs: the relative state of the frame itself factorises out [8] and is therefore sometimes omitted [2]. [^11] Note that while this is taken as a physically motivated assumption in [8], it is argued for in [22] by specifying the conditions under which a frame satisfies a localisability property. Since this always holds for what we refer to as perfect reference frames (see Sec. IV), this can be seen as additionally corroborating this assumption. Now, if the two-particle system is described, for example, by Newtonian gravity, the overall configuration of the two systems is translation-invariant. For simplicity, we exclude the rotational symmetry of the theory and consider only the invariance under translations in one dimension. As a consequence, any two models that are related by a rigid translation live on the same orbit. This is the case for models with the same relative distance $d_{MP}=x_{M}-x_{P}$. Thus in this example, each orbit can be labelled by the relative distance between the two systems $P$ and $M$. As the (signed) relative distance between $M$ and $P$ differs across the branches in our setup, we are dealing with a superposition of two models that are physically distinguishable and therefore live on different orbits in the space of models (see Fig. 6). [^12] Note that we would consider superpositions of models living on the same orbit as “false” superpositions. In the present context, this would correspond to a superposition of two configurations with the same relative distance between $M$ and $P$ but distinct centre of mass position.
@@
Consider the following example, which was worked out in [10], of a three-particle system, whose positions are denoted by $q_{1},q_{2},q_{3}$. When applying a standard, model-independent translation, the position of each system is shifted by the same fixed amount: $q_{i}\to q_{i}+a$, $i=1,2,3$. For the model-dependent translation, on the other hand, the amount by which the positions are shifted can depend on the model $\varphi=(q_{1},q_{2},q_{3})$ and thus on the configuration of the other particles. This is the case, for example, when changing from the reference frame of Particle 1 into the reference frame of Particle 2, even at the classical level. In the former, what we mean by “the position of Particle 3” is the “the position of Particle 3 relative to Particle 1”, that is $Q_{3}^{(1)}=q_{3}-q_{1}$. Upon changing into the reference frame of Particle 2, on the other hand, “the position of Particle 3” is now $Q_{3}^{(2)}=Q_{3}^{(1)}+q_{1}-q_{2}=q_{3}-q_{2}$, namely “the position of Particle 3 relative to Particle 2”. In this sense, this transformation “mixes” the degrees of freedom of the different subsystems and thereby changes the notion of subsystem. While this change of subsystem structure is a well-understood feature at the classical level, it develops new implications at the quantum level, where it is reflected in a refactorisation of the Hilbert space and thereby the frame-dependence of entanglement.